# REA-RL: Reflection-Aware Online Reinforcement Learning for Efficient Large Reasoning Models

### v1

@@ Appendix A Prompts
  • The Prompts for Detecting Overthinking: These prompts are used in §3.1. We first segment the response into chunks, specifically by splitting it at every two newline characters. To prevent a single formula from being split across multiple chunks, we merge chunks until the combined chunk ends with a sentence-ending punctuation mark or exceeds 128 tokens. To prevent the model from generating excessively long labeling strings, we input a maximum total length of 1k tokens at a time. Subsequently, we first feed all the chunks into the first prompt for an initial detection. Each chunk subsequently labeled as "Right Result" is then fed into the second prompt for secondary labeling. "With / without Gold Answer" indicates whether the gold answer is included as input.
  • The SFT Prompt for Reflection Model Training: These prompts constitute the SFT data we constructed in §4.2 to train a 7B reflection model. We employ simpler definitions and non-chain-of-thought output to ensure both effectiveness and efficiency. The gold answer is not used as input to ensure usability during training and evaluation.
@@
The First Prompt for Detecting Overthinking without Gold Answer **Question:** {Question} **Response:** {All Chunked Responses} You are provided with a math Question and a model-generated Response. The response is divided into {N} parts. For each part, analyze it and classify it based on its relationship to the provided context. For each part, assign one of the following labels: - Reasoning: The part represents the reasoning process that leads to the answer. - Right Result: The part is the answer provided by the model, where the model may provide the answer in the middle of its response, and the answer align with the Gold Answer. - Wrong Result: Same as Right Result, but the answer do not align with the Gold Answer. For each of the {N} parts, please reply in format: [1]. Think: [Explanation for label choice] Label: Reasoning/Right Result/Wrong Result [2]. Think: [Explanation for label choice] Label: Reasoning/Right Result/Wrong Result
  • The Prompt for Math Evaluation: We follow Yang et al. [42] to design the prompts for math evaluation. When forcing the model to complete generation and provide the final answer, we append “</think> **Final Answer:**” after the thought process and allow the model to continue generating up to 1k tokens. During training, to ensure efficiency, we reduce the budget to 256 tokens.
The Second Prompt for Detecting Overthinking with Gold Answer **Question:** {Question} **Gold Answer:** {Answer} **Response:** {One Chunked Response} Evaluate whether the model correctly answered the question. As long as the model provides the correct result, it counts as correct, regardless of format or wording. The response I provided is part of the complete response, so thereś no need to include the entire reasoning process. Please judge only if the model has provided the correct answer up to this point. Please reason step by step first after "Reasoning:", then answer only with Yes or No after "Answer:".
The Second Prompt for Detecting Overthinking without Gold Answer **Question:** {Question} **Response:** {One Chunked Response} Evaluate whether the model has already answered the question. The response I provided is part of the complete response, so thereś no need to include the entire reasoning process. Please judge only if the model has provided the answer up to this point. Please reason step by step first after "Reasoning:", then answer only with Yes or No after "Answer:".
@@ B.1 Case Study for REA-RL
To illustrate how our method works, we provide a case study demonstrating the workflow of online data generation for online RL, containing both parallel sampling and sequential revision, as shown in Figure 4. Specifically, we first perform parallel sampling, obtaining the yellow and red parts. Here, the red part represents the overthinking section, which is detected by the reflection model. Subsequently, we perform sequential revision by removing the red part and allowing the policy model to finish the response with the blue part. In detail, we force the termination of the thinking process by adding a “</think>” token and enforce answer generation by adding “**Final Answer:**” to avoid further redundant reasoning. Based on this process, the blue part is generally much shorter than the red part, thus providing cases with less overthinking online.
@@
Figure 4: Case Study for REA-RL of online data generation. We illustrate how the parallel sampling and sequential revision parts work. Specifically, the yellow, red, and blue parts in this figure correspond to the tokens of the same colors in Figure 2. Since the yellow parts in both completion and revision are identical, they are shown only once in this figure.
B.2 Ablation Study of Reward Refinement
@@
Across all experiments, we observe an average performance gain of 5% compared to the scenarios without the reflection reward. However, this improvement comes at the cost of an average 16% reduction in the truncation ratio. Nevertheless, we believe that maintaining performance is a more critical objective than shortening the response. Only with the addition of the reflection reward can we achieve a reduction in model response length without sacrificing model performance, and we anticipate further shortening of the response with continued training, as illustrated in Figure 3. Conversely, continuing training with only the length reward may lead to a further decline in accuracy.
@@
Method GSM8K Math500 Gaokao23 Amc23 Olympiad Aime24 Average
@@
Budget: 8k Original 91.58 1001312 88.40 1003389 77.40 1003270 77.50 1004746 51.70 1005467 40.83 1006815 71.24 1004166
GRPO $\text{R}_{\text{Len}}$ 85.97 32.01 87.60 59.04 70.91 58.84 85.62 71.43 55.56 76.59 45.42 87.42 71.85 64.22
GRPO $\text{R}_{\text{RLen}}$ 86.05 29.42 88.20 50.19 68.83 51.10 86.88 63.27 53.93 71.67 48.75 84.92 72.11 58.43
GRPO $\text{R}_{\text{Len+Reflect}}$ 92.95 89.63 91.60 81.74 81.30 84.68 83.12 87.00 55.11 88.38 44.58 95.52 74.78 87.83
 w/ $D_{0.1}$ 92.34 80.26 91.60 76.54 79.22 78.87 84.06 82.64 53.78 87.91 40.00 93.25 73.50 83.25
 w/ $D_{0.4}$ 93.25 108.46 88.00 93.24 80.52 96.51 81.88 92.82 52.59 95.43 37.50 100.18 72.29 97.77
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 67.07 91.40 69.52 80.52 72.54 85.62 75.64 53.93 83.04 47.92 91.37 75.35 76.53
Budget: 16k Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 60.15 1007549 48.33 10010460 77.01 1005478
GRPO $\text{R}_{\text{Len}}$ 85.97 31.25 88.40 56.43 72.21 55.88 87.81 65.65 59.56 69.20 50.00 76.96 73.99 59.23
GRPO $\text{R}_{\text{RLen}}$ 86.05 28.72 89.20 46.62 69.35 47.03 88.12 56.59 56.74 64.47 54.17 75.40 73.94 53.14
GRPO $\text{R}_{\text{Len+Reflect}}$ 92.95 88.10 93.20 79.58 82.86 79.39 87.50 83.30 59.70 83.32 51.67 89.54 77.98 83.87
 w/ $D_{0.1}$ 92.34 78.65 93.20 73.67 81.56 77.31 89.06 80.62 61.33 82.50 47.92 91.85 77.57 80.77
 w/ $D_{0.4}$ 93.40 107.59 91.60 93.68 83.90 92.26 89.69 92.47 60.15 93.16 46.25 101.19 77.50 96.72
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 65.48 92.80 66.71 81.82 68.27 88.75 72.17 58.81 77.37 54.58 86.21 78.25 72.70
B.3 Overthinking Analysis After Training
@@
To verify that our method alleviates the overthinking issue, we employ the same evaluation approach as in §3 on our trained model, which uses an LLM to identify overthinking tokens and remove them with model reflection. The results are presented in Table 6. Compared to directly truncating the response from the original model, our approach exhibits a considerably lower truncation ratio for overthinking tokens, especially for methods employing our reflection model for training. This demonstrates that REA-RL effectively mitigates the issue of overthinking.
@@
Method GSM8K Math500 Gaokao23 Amc23 Olympiad Aime24 Average
[…]
Budget: 16k Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 60.15 1007549 48.33 10010460 77.01 1005478
Model Reflect 89.46 62.43 93.20 71.85 80.52 70.41 89.06 79.95 60.74 83.06 50.83 93.59 77.30 76.88
 + Gold 88.78 54.09 92.00 57.80 79.48 59.71 87.19 67.71 60.15 73.48 51.67 92.31 76.54 67.52
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 100880 92.80 1002597 81.82 1002584 88.75 1004215 58.81 1005841 54.58 1009018 78.25 1004189
Model Reflect 92.42 72.73 93.00 78.17 80.26 75.46 87.19 84.01 57.93 86.77 54.58 94.72 77.56 81.98
 + Gold 91.74 68.75 91.60 69.77 79.74 70.51 85.94 77.27 57.19 79.27 54.58 93.70 76.80 76.54
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 100496 92.40 1001756 79.74 1001798 88.12 1003273 60.74 1004795 47.92 1007931 76.36 1003342
Model Reflect 88.02 90.73 91.40 92.54 79.22 92.99 88.75 93.92 59.41 91.60 45.42 95.51 75.37 92.88
 + Gold 87.72 88.91 89.80 84.85 78.96 88.65 86.56 90.53 58.96 83.96 46.25 95.21 74.71 88.68
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 100700 91.00 1002366 82.08 1002114 89.38 1003957 59.11 1005519 51.25 1008482 77.13 1003856
Model Reflect 90.30 86.29 91.60 88.08 80.78 88.69 88.75 90.90 60.59 90.20 50.42 96.24 77.07 90.07
 + Gold 89.46 80.86 90.40 80.85 79.74 82.64 87.19 88.10 59.41 82.68 50.83 95.48 76.17 85.10

### v2

@@ Appendix A Prompts
[…]
  • The SFT Prompt for Reflection Model Training: These prompts constitute the SFT data in §4.2 to train a 7B reflection model. We employ simpler definitions and non-chain-of-thought output to ensure both effectiveness and efficiency. The gold answer is not used as input to ensure usability during training and evaluation.
@@
The First Prompt for Detecting Overthinking without Gold Answer **Question:** {Question} **Response:** {All Chunked Responses} You are provided with a math Question and a model-generated Response. The response is divided into {N} parts. For each part, analyze it and classify it based on its relationship to the provided context. For each part, assign one of the following labels: - Reasoning: The part represents the reasoning process that leads to the answer. - Right Result: The part is the answer provided by the model, where the model may provide the answer in the middle of its response, and the answer aligns with the Gold Answer. - Wrong Result: Same as Right Result, but the answer does not align with the Gold Answer. For each of the {N} parts, please reply in format: [1]. Think: [Explanation for label choice] Label: Reasoning/Right Result/Wrong Result [2]. Think: [Explanation for label choice] Label: Reasoning/Right Result/Wrong Result
  • The Prompt for Math Evaluation: We follow Yang et al. (2024b) to design the prompts for math evaluation. For NoThink, we append “Okay, I think I have finished thinking. ¡/think¿” after the prompt to force LRMs to skip reasoning. When forcing the model to complete generation and provide the final answer, we append “¡/think¿ **Final Answer:**” after the thought process and allow the model to continue generating up to 1k tokens. During training, to ensure efficiency, we reduce the budget to 256 tokens.
The Second Prompt for Detecting Overthinking with Gold Answer **Question:** {Question} **Gold Answer:** {Answer} **Response:** {One Chunked Response} Evaluate whether the model correctly answered the question. As long as the model provides the correct result, it counts as correct, regardless of format or wording. The response I provided is part of the complete response, so thereś no need to include the entire reasoning process. Please judge only if the model has provided the correct answer up to this point. Please reason step by step first after ”Reasoning:”, then answer only with Yes or No after ”Answer:”.
The Second Prompt for Detecting Overthinking without Gold Answer **Question:** {Question} **Response:** {One Chunked Response} Evaluate whether the model has already answered the question. The response I provided is part of the complete response, so thereś no need to include the entire reasoning process. Please judge only if the model has provided the answer up to this point. Please reason step by step first after ”Reasoning:”, then answer only with Yes or No after ”Answer:”.
@@ B.1 Case Study for REA-RL
Data filtering (Deng et al., 2023; 2025; Liu et al., 2024b; Rao et al., 2025; Wang et al., 2025) and data synthesis (Deng et al., 2024; Ke et al., 2025; Yuan et al., 2025) are important means of enhancing model capabilities. However, because these techniques are relatively time-consuming, comparatively few studies have explored deploying them in an online setting. To further illustrate how our method works, we provide a case study demonstrating the workflow of online data generation for online RL, containing both parallel sampling and sequential revision, as shown in Figure 4. Specifically, we first perform parallel sampling, obtaining the yellow and red parts. Here, the red part represents the overthinking section, which is detected by the reflection model. Subsequently, we perform sequential revision by removing the red part and allowing the policy model to finish the response with the blue part. In detail, we force the termination of the thinking process by adding a “¡/think¿” token and enforce the generation of the answer by adding “**Final Answer:**” to avoid further redundant reasoning. Based on this process, the blue part is generally much shorter than the red part, thus providing positive cases with less overthinking online.
@@
Figure 4: Case study of online data generation in REA-RL. We illustrate how the parallel sampling and sequential revision parts work. Specifically, the yellow, red, and blue parts in this figure correspond to the tokens of the same colors in Figure 2. Since the yellow parts in both completion and revision are identical, they are shown only once in this figure.
B.2 Additional Experimental Setup
Common training configuration.
Unless explicitly stated otherwise, we maintain the following experimental settings. We use the DeepScaleR 40k dataset (Luo et al., 2025b) as the training data and DeepSeek-R1-Distill-Qwen-7B (R1-7B) as the base model. We only retain questions for which the model can provide at least one correct answer within 4 samples as the training data, resulting in 30k questions. We train for one epoch for GRPO-based methods and the same number of steps for all other methods and baselines. Each batch contains 16 different questions and 4 different paths generated for each question, with a maximum generation length of 8k. We use a learning rate of $2\times 10^{-5}$, integrating low-rank adaptation (Hu et al., 2022) with all attention and MLP parameters, setting the LoRA rank $r$ to 16 and alpha to 32. During the data generation process, we follow R1 to use a temperature of 0.6 and a top_p of 0.95. Finally, following Yu et al. (2025), when all generated trajectories yield incorrect answers, we skip these instances to avoid unintended optimization.
$\text{M}_{\text{Reflect}}$ R1-1.5B DeepSeek-7B Average
[…]
Original 30.30 1004844 49.32 1003834 39.81 1004339
Qwen2.5-7B-Instruct-Weak 30.62 100.72 50.42 99.40 40.52 100.06
Qwen3-4B-Weak 30.48 101.96 51.60 99.48 41.04 100.72
Qwen2.5-7B-Instruct-Normal 31.40 95.11 50.12 92.96 40.76 94.04
Qwen3-4B-Normal 31.40 96.26 50.42 94.31 40.91 95.29
Qwen2.5-7B-Instruct-Strong 30.60 89.91 48.44 87.14 39.52 88.53
Qwen3-4B-Strong 31.16 89.88 48.70 87.48 39.93 88.68
Table 5: General-QA reflection model evaluation on MMLU-Pro using the truncation-and-force-answer protocol. Other abbreviations are defined in Table 3.
Method MMLU MMLU-Pro SuperGPQA Average
[…]
Budget: 8k
Original 40.42 1001598 27.44 1003632 13.52 1004441 27.13 1003214
GRPO 40.36 74.55 28.10 85.77 11.98 98.13 26.81 86.15
GRPO + $\text{M}_{\text{Reflect}}$ 40.94 73.85 27.88 89.98 13.56 93.22 27.46 85.68
Budget: 16k
Original 40.98 1001871 30.30 1004844 16.66 1006337 29.31 1004351
GRPO 40.58 70.82 30.24 84.56 15.36 104.92 28.73 86.77
GRPO + $\text{M}_{\text{reflect}}$ 41.30 74.77 30.48 89.66 16.86 91.68 29.55 85.37
Table 6: General-QA policy training results on general QA. Abbreviations are defined in Table 2.
Method GSM8K Math500 GaoKao23 Amc23 Aime Average
[…]
Budget: 8k
Original 75.51 100615 67.80 1004137 65.97 1003336 52.81 1005412 18.33 1007581 56.08 1004216
GRPO 74.83 58.54 77.40 52.28 63.64 65.14 60.62 72.60 22.92 86.26 59.88 66.96
GRPO + $\text{M}_{\text{Reflect}}$ 72.33 60.81 80.00 55.14 64.68 67.51 65.31 73.50 24.58 88.83 61.38 69.16
Budget: 16k
Original 75.59 100622 72.00 1005652 69.61 1004327 57.81 1008283 21.25 10013358 59.25 1006448
GRPO 74.83 58.84 78.60 46.16 65.71 59.88 63.75 60.69 24.58 79.45 61.49 61.00
GRPO + $\text{M}_{\text{reflect}}$ 72.33 60.13 81.60 48.73 68.05 59.51 69.06 62.77 28.75 75.58 63.96 61.34
Table 7: Additional math-domain results on R1-1.5B. Abbreviations are defined in Table 2.
B.3 Generalization Results on Broader Domains and Model Scales
To address the concern that our main experiments focus on the math domain, we add experiments on general QA reasoning and smaller LRMs following prior evaluation setups.
Models and reflection model selection.
We intentionally use smaller policy models to verify generality in model scales. Specifically, we report results on DeepSeek-R1-Distill-Qwen-1.5B (R1-1.5B). For the general QA setting, we also train a new reflection model: we generate SFT data with R1-1.5B as the trajectory generator and Qwen3-32B (Yang et al., 2025) as the reviser, mix these QA-domain data with the math-domain data, and train reflection models based on Qwen2.5-7B-Instruct and Qwen3-4B.
General QA domain dataset.
Concretely, we evaluate on MMLU (Hendrycks et al., 2021a), MMLU-Pro (Wang et al., 2024), and SuperGPQA (Team et al., 2025), covering different difficulty levels. Based on the dataset sizes, we select 20k examples from MMLU, 5k from MMLU-Pro, and 15k from SuperGPQA to construct a 40k mixed training set, and we use 5k examples from each dataset as test sets. Since some QA questions do not require deep reasoning, we first filter out the 15% of questions with the shortest paths (generated by R1-1.5B), and then sample the required number of examples. Finally, for the training set, we remove questions that R1-1.5B cannot solve within 4 paths, resulting in a 30k-example training set.
Policy training setup.
[…]
Reflection model evaluation results on general QA.
Under the same truncation-and-force-answer evaluation protocol (tested on MMLU-Pro), the trained reflection models can effectively truncate trajectories with comparable performance across 7B and 4B reflection models, as shown in Table 5. Across reflection model sizes, we observe a consistent trade-off controlled by the truncation strength: the Normal setting noticeably reduces TR while maintaining or slightly improving average accuracy, whereas Strong further reduces TR but can introduce a larger accuracy drop. Overall, these results indicate that the reflection model transfers to the QA domain under the same evaluation protocol, and that smaller reflection models (4B) can closely match the 7B reflection model in both accuracy and truncation behavior. Therefore, we use the 4B model as the reflection model for general QA in the following experiments.
Policy training results on general QA and math.
Table 6 reports policy training results on general QA with R1-1.5B. Under both 8k and 16k budgets, introducing the reflection model yields stable improvements in average accuracy while maintaining a comparable average token ratio. Table 7 further reports additional experiments in the math domain on R1-1.5B, using the same training and generation configuration as in the main experiments: under both budgets, our approach improves average accuracy compared to GRPO while achieving comparable average token ratios. Taken together, these results demonstrate that the approach generalizes to smaller models across both math and general QA domains.
B.4 Ablation Study of Reward Refinement
Method GSM8K Math500 Gaokao23 Amc23 Aime24 Average
[…]
Budget: 8k Original 91.58 1001312 88.40 1003389 77.40 1003270 77.50 1004746 40.83 1006815 75.14 1003906
GRPO $\text{R}_{\text{Len}}$ 85.97 32.01 87.60 59.04 70.91 58.84 85.62 71.43 45.42 87.42 75.10 61.75
GRPO $\text{R}_{\text{RLen}}$ 86.05 29.42 88.20 50.19 68.83 51.10 86.88 63.27 48.75 84.92 75.74 55.78
GRPO $\text{R}_{\text{Len+Reflect}}$ 92.95 89.63 91.60 81.74 81.30 84.68 83.12 87.00 44.58 95.52 78.71 87.71
  w/ $D_{0.1}$ 92.34 80.26 91.60 76.54 79.22 78.87 84.06 82.64 40.00 93.25 77.44 82.31
  w/ $D_{0.4}$ 93.25 108.46 88.00 93.24 80.52 96.51 81.88 92.82 37.50 100.18 76.23 98.24
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 67.07 91.40 69.52 80.52 72.54 85.62 75.64 47.92 91.37 79.64 75.23
Budget: 16k Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 48.33 10010460 80.39 1005064
GRPO $\text{R}_{\text{Len}}$ 85.97 31.25 88.40 56.43 72.21 55.88 87.81 65.65 50.00 76.96 76.88 57.23
GRPO $\text{R}_{\text{RLen}}$ 86.05 28.72 89.20 46.62 69.35 47.03 88.12 56.59 54.17 75.40 77.38 50.87
GRPO $\text{R}_{\text{Len+Reflect}}$ 92.95 88.10 93.20 79.58 82.86 79.39 87.50 83.30 51.67 89.54 81.64 83.98
  w/ $D_{0.1}$ 92.34 78.65 93.20 73.67 81.56 77.31 89.06 80.62 47.92 91.85 80.82 80.42
  w/ $D_{0.4}$ 93.40 107.59 91.60 93.68 83.90 92.26 89.69 92.47 46.25 101.19 80.97 97.44
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 65.48 92.80 66.71 81.82 68.27 88.75 72.17 54.58 86.21 82.13 71.77
@@ B.4 Ablation Study of Reward Refinement
Across all experiments, we observe an average performance gain of 4.26 compared to the scenarios without the reflection reward. However, this improvement comes at the cost of an average 23.26 reduction in the truncation ratio. Nevertheless, we believe that maintaining performance is a more critical objective than shortening the response. Only with the addition of the reflection reward can we achieve a reduction in model response length without sacrificing model performance, and we anticipate further shortening of the response with continued training, as illustrated in Figure 3. Conversely, continuing training with only the length reward may lead to a further decline in accuracy.
@@
B.5 Token Cost Analysis and Training-Time Overhead
Token consumption and cost.
We measure token usage per path during training in our main experiments. For parallel sampling, the average input length is 94 tokens and the average output length is 3605 tokens. For sequential revision, the input is 6678 tokens and the output is 319 tokens in total. Given that input tokens are typically at most 1/5 the price of output tokens, converting output tokens into input tokens can substantially reduce cost under large-scale generation. Using publicly available API pricing as an example, for GPT-5, the input price is \$1.25 per 1M tokens (\$0.125 for cached input), and the output price is \$10.00 per 1M tokens. Under this pricing, parallel sampling for 1k examples costs \$36.17, whereas sequential revision costs \$7.78.
Further, we measure small-scale vLLM inference time on a per-batch basis, which aligns with our training setup. On a single A800 GPU, generating one batch with 4 paths takes 158s; sequential revision takes 47s for labeling plus 2s for answer generation (49s total), which is less than one third of the traditional rollout. In our online pipeline, these stages run sequentially per batch (158 + 49 = 207s), which can be slower than pure parallel sampling due to reduced parallelism, i.e., sampling 8 paths in parallel takes 192s. Therefore, in practice we spend more time on training. In our setup on 3 A800 80G GPUs, sampling 4 paths takes 80h, sampling 8 paths takes 110h, and incorporating sequential revision takes 120h. However, at larger generation scale this loss of parallelism is mitigated, bringing the final cost closer to the price ratio estimated by the above API calls.
Heuristic truncation at reflection tokens.
Although introducing a reflection model incurs additional latency, we find it necessary, as heuristic truncation is unlikely to achieve consistently strong results. We add a baseline that truncates the reasoning trace at the first (or fifth) sentence that begins with a reflection token, and then forces final-answer generation. Results are shown in Table 9. Truncating at the first occurrence is too aggressive and causes a large accuracy drop, because reflection can occur during intermediate reasoning rather than only after finishing. Truncating at the fifth occurrence improves results on easier datasets but still yields substantial drops on harder datasets, indicating that the optimal truncation strength is difficulty-dependent and hard to tune; this supports the necessity of a learned reflection model.
Method GSM8K Math500 Gaokao23 Amc23 Aime24 Average
@@ B.5 Token Cost Analysis and Training-Time Overhead
Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 48.33 10010460 80.39 1005064
First 69.37 29.84 49.80 14.98 43.38 15.67 29.06 11.92 2.08 5.87 38.74 15.66
Fifth 89.69 68.01 70.80 40.41 60.52 38.57 55.00 29.76 16.25 15.50 58.45 38.45
32B Revise 89.46 62.43 93.20 71.85 80.52 70.41 89.06 79.95 50.83 93.59 80.61 75.65
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 52.08 91.00 60.78 82.08 55.85 89.38 67.76 51.25 81.09 80.74 63.51
Table 9: Heuristic truncation baselines vs. learned truncation model. Other abbreviations are defined in Table 3.
B.6 Overthinking Analysis After Training
Method GSM8K Math500 Gaokao23 Amc23 Aime24 Average
[…]
Budget: 16k Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 48.33 10010460 80.39 1005064
Model Reflect 89.46 62.43 93.20 71.85 80.52 70.41 89.06 79.95 50.83 93.59 80.61 75.65
  + Gold 88.78 54.09 92.00 57.80 79.48 59.71 87.19 67.71 51.67 92.31 79.82 66.32
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 100880 92.80 1002597 81.82 1002584 88.75 1004215 54.58 1009018 82.13 1003859
Model Reflect 92.42 72.73 93.00 78.17 80.26 75.46 87.19 84.01 54.58 94.72 81.49 81.02
  + Gold 91.74 68.75 91.60 69.77 79.74 70.51 85.94 77.27 54.58 93.70 80.72 76.00
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 100496 92.40 1001756 79.74 1001798 88.12 1003273 47.92 1007931 79.48 1003051
Model Reflect 88.02 90.73 91.40 92.54 79.22 92.99 88.75 93.92 45.42 95.51 78.56 93.14
  + Gold 87.72 88.91 89.80 84.85 78.96 88.65 86.56 90.53 46.25 95.21 77.86 89.63
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 100700 91.00 1002366 82.08 1002114 89.38 1003957 51.25 1008482 80.74 1003524
Model Reflect 90.30 86.29 91.60 88.08 80.78 88.69 88.75 90.90 50.42 96.24 80.37 90.04
  + Gold 89.46 80.86 90.40 80.85 79.74 82.64 87.19 88.10 50.83 95.48 79.52 85.59
@@ B.6 Overthinking Analysis After Training
To verify that our method alleviates the overthinking issue, we employ the same evaluation approach as in §3, which uses an LLM to identify overthinking tokens and remove them with model reflection. The results are presented in Table 10. Compared to directly truncating the response from the original model, our approach exhibits a considerably lower truncation ratio for overthinking tokens, especially for methods employing our reflection model for training. This demonstrates that REA-RL effectively mitigates the issue of overthinking.
@@
B.7 The Use of Large Language Models
As a general-purpose assist tool, large language models are used only to aid writing (e.g., grammar, clarity, and phrasing). They are not used to design methods, generate ideas, run experiments, or create results. All technical content and conclusions are written and verified by the authors.